Mountain Dental Services is a specialized dental practice whose only service is filling cavities. Mountain has recorded the following for the past nine months:

Month Number of Cavities Filled Total Cost
January 400 $5,350
February 325 5,050
March 675 6,600
April 700 6,300
May 425 5,450
June 500 5,850
July 550 6,150
August 300 5,200
September 450 5,500

Required:

1. Use the high-low method to estimate total fixed cost and variable cost per cavity filled.

2. Using these estimates, calculate Mountain’s total cost for filling 800 cavities.
Answer #1

Variable cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (6300-5200)/(700-300) = 2.75

Fixed cost = 6300 - (700*2.75) = 4375

Fixed cost 4375
Variable cost per unit 2.75

Total estimated cost = (2.75*800) + 4375 = 6575